OpenStudy (anonymous):

the ratio of boys to girls in the auditorium was 4 to 5.if there were 40 boys in the auditiorium how many girls were there?

OpenStudy (anonymous):

1 book costs 1.95 so the cost of 12 books will be 12 times 1.95 i.e 12*1.95

OpenStudy (anonymous):

4/5=40/x By cross multiplication, 4x=200. Therefore there are 50 girls.

OpenStudy (anonymous):

oh sorry i commented on the wrong post i guess